Two large non-conducting plates of surface ares A - 0 25 m2 cay equal but opposate charges Q-75uc wWhat i the energy demsty of the electric teld between the h plates? O 25 1015m3 0 34 107 Wm3 O 5.1 1015 m3 O 1.7 107 Jm3
Answer #1

energy density = 1/2 * \epsilono * E^2

= 1/2 * \epsilono * (Q/A\epsilono ) ^2

= Q^2 / 2A^2\epsilono

= (75*10^-6)^2 / 2*0.25^2*8.85*10^-12

= 5084 J/m^3